6. Allen Lumber Company had earnings after taxes of $600,000 in the year 2009 with 400,000 shares outstanding on December 31, 2009. On January 1, 2010, the firm issued 34,000 new shares. Because of the proceeds from these new shares and other operating improvements, 2010 earnings after taxes were 25 percent higher than in 2009. Earnings per share for the year 2010 were (Round your answer to 2 decimal places.) O $1.88 O $1.51 O $1.73 O $1.38
Answer #1

Calculation if earning per share for 2010:

Earning per share= earning after tax/number of shares

Earning after tax= 600000*1.25= 750000

Number of shares= 400000+34000= 434000

Earning per share= 750000/434000= 1.73

So correct answer is $1.73